OpenStudy (anonymous):

Solve: 2(x + 1) = −4

OpenStudy (anonymous):

distribute the 2 2(x+1)= 2*x+1*2 then solve for x

OpenStudy (anonymous):

2x+2=-4 -2 -2 2x=-6 2 2 anwser=-3

OpenStudy (anonymous):

Thanks Guys! This is one of my lazy days lol

OpenStudy (anonymous):

lol ikr !

OpenStudy (anonymous):

Yeah!